A pump is driven by an electric motor running at 1145 rpm and the speed of the compressor is 920 rpm. If the diameter of the motor pulley is 160 mm, calculate the diameter of the pump pulley


1
Expert's answer
2021-02-15T05:42:54-0500

Speed of motor = 1145 rpm

Diameter of the motor pulley = 160mm


Speed of the compressor = 920rpm

Diameter of the compressor = xmm


"\dfrac{1145}{160} = \dfrac{x}{920}"


"x = \dfrac{1145×920}{160}" = 6584mm